OpenStudy (anonymous):

Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2??

OpenStudy (lgbasallote):

hint!!  does that help?

OpenStudy (anonymous):

yah, kinda. It would help if I seen work. srry.

OpenStudy (lgbasallote):

hmm okay let me present another hint...the perpendicular line would be a horizontal line correct?

OpenStudy (anonymous):

uh-huh.

OpenStudy (anonymous):

the formula for this is y=mx+b?

OpenStudy (lgbasallote):

actually...you cant use y= mx + b here because there is no y in x = 2 there can be a formula here but it's complicated to explain so im trying it graphically

OpenStudy (lgbasallote):

would you like to hear the formula though?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yah.

OpenStudy (lgbasallote):

okay...so the line is x = 2 as i showed you..the line is vertical. it is a rule in this thing that if the graph is vertical then it has 0 slope if it is horizontal then undefined slope do you get that?

OpenStudy (anonymous):

yah, but dnt u mean -2?

OpenStudy (lgbasallote):

why would you think it's -2?

OpenStudy (anonymous):

you u mean the (2,1) <--that's in here, that's -2?

OpenStudy (lgbasallote):

no..im talking about the line x = 2...the one it's saying the line is perpendicular to?

OpenStudy (anonymous):

oh, okay.

OpenStudy (agent47):

y=1

OpenStudy (agent47):

x=-2 has a vertical slope so you need a horizontal slope, or slope 0

OpenStudy (lgbasallote):

i made a mistake in my earlier comment slope of vertical is undefined then slope of horizontal is 0

OpenStudy (agent47):

y-y1=0(x-x1) y-1=0 y=1

OpenStudy (anonymous):

thank you! 2 the both of u :)